### v1

# NeurIPS 2025 Workshop on Time Series for Health

Abstract
Reinforcement learning (RL) applications in Clinical Decision Support Systems (CDSS) frequently encounter skepticism from practitioners regarding inoperable dosing decisions. We address this challenge with an end-to-end approach for learning optimal drug dosing and control policies for dual vasopressor administration in intensive care unit (ICU) patients with septic shock. For realistic drug dosing, we apply action space design that accommodates discrete, continuous, and directional dosing strategies in a system that combines offline conservative Q-learning with a novel recurrent modeling in a replay buffer to capture temporal dependencies in ICU time-series data. Our comparative analysis of norepinephrine dosing strategies across different action space formulations reveals that the designed action spaces improve interpretability and facilitate clinical adoption while preserving efficacy. Empirical results [^1] Code for this paper is publicly available at https://github.com/wzoustanford/ucsf_rl/ on eICU and MIMIC demonstrate that action space design profoundly influences learned behavioral policies. The proposed methods achieve improved patient outcomes of over 15% in survival improvement probability, while aligning with established clinical protocols.
1 Introduction
Reinforcement learning recommendations are promising to serve as clinician support, but when a seasoned clinician encounters a proposed dosing adjustment as small as $0.01$ mg, the model’s credibility erodes. This disconnect illuminates a need to bridge the gap between theoretically optimal RL policies and recommendations that are clinically actionable, interpretable, and trusted.
Clinical interventions in Intensive Care Units (ICU’s) are inherently heterogeneous, spanning discrete choices and continuous adjustments jayaraman2024primer ; eghbali2021patient ; qiu2022latent ; tosca2024model . This diversity necessitates action space formulations that can flexibly capture the full spectrum of clinical practice while remaining tractable for RL algorithms. Without careful action space design, RL policies risk generating unrealistic or impractical recommendations that clinicians are unlikely to adopt. More specifically, in a Clinical Decision Support System (CDSS) hunt1998effects ; garg2005effects for septic shock, vasopressor dosing requires dynamic control across multiple phases—initiation, titration, and weaning. Current clinical practices adhere to international standards of applying norepinephrine as the first line vasopressor and vasopressin as the second-line vasopressor. While reinforcement learning offers a natural framework for this sequential decision-making problem, existing work kalimouttou2025optimal ; komorowski2018artificial has predominantly focused on single vasopressor control and early treatment phases, overlooking the reality of titration during the full length of therapy.
To address these challenges, we investigate the patterns of both vasopressors, especially the impact of norepinephrine dosing patterns with effective action space design for clinical drug dosing towards a realistic CDSS system. Concretely, we adopt the offline settings of the Deep Q-learning (DQN) mnih2015human with recurrent replay kapturowski2018recurrent , and conservative Q-learning (CQL) kumar2020conservative . Integrating them together as a potential CDSS system, we validate the action space design for dosing in treatment trajectories.
We find that discrete and directional-discrete action spaces are not only more interpretable and aligned with clinician expectations, but also improve survival probability by over 15% compared to continuous dosing. Their sparsity and higher dimensionality, combined with recurrent and replay frameworks on short sequences, enable more generalizable models. These results show the potential of deep reinforcement learning and action-space design for building effective CDSS systems.
2 Prior Work
Reinforcement learning techniques have been applied in several independent studies for treating sepsis shock. For instance,  saria2018individualized proposed individualized treatment for sepsis using RL, and komorowski2018artificial provided an earlier comprehensive study of sepsis treatment with publicly available datasets. More recently, jayaraman2024primer studied ICU data and RL in medicine. kalimouttou2025optimal provided a focused study of effective dosing action space design in an end-to-end RL system for ICU sepsis treatment.
For action space design in RL, earlier work addressed factored tavakoli2018action , hierarchical kulkarni2016hierarchical , continuous action spaces tessler2019action ; delalleau2019discrete kapturowski2018recurrent addressed the importance of recurrent sequence learning in replays for application in Atari games, and  schrittwieser2020mastering progressed recurrent learning into model-based RL applications. Offline RL methods such as conservative Q-learning kumar2020conservative , implicit Q-learning kostrikov2021offline and advantage weighted Actor-Critic nair2020awac applied different regularization and policy constraints. Model-based offline RL approaches kidambi2020morel focused on modeling transition dynamics. In the clinical literature, eghbali2021patient applied RL algorithm to sedation dosing in ICUs. A collection of dosing techniques was given in  tosca2024model , focusing on referencing cancer and chemotherapy treatments.  qiu2022latent applied a latent constrained batch algorithm to address the risks of incorrect dosing heparin in ICU. However, the use of RL to systematically design and evaluate vasopressor dosing strategies for sepsis shock remains under-explored.
3 Algorithms
We propose an offline Q-learning framework with recurrent experience replay, focusing on effective dosing action-spaces with conservatism objective from conservative Q-learning.
3.1 Vasopressor Control Problem Formulation
We formulate the vasopressor control problem with the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space of vasopressor decisions, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the state transition dynamics, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward, and $\gamma\in[0,1]$ is the discount factor. The state $s_{t}\in\mathcal{S}\subseteq\mathbb{R}^{d}$ comprises patient information, physiological measurements [^2] The measurements include mean blood pressure (MBP), lactate levels, blood urea nitrogen, serum creatinine, (total) fluid, urine of the hour, corticosteroid and Sequential Organ Failure Assessment (SOFA) score. and organ support [^3] The support includes mechanical ventilation, renal replacement therapy.. Action space $\mathcal{A}$ contains a binary decision for vasopressin ($vp_{1}$) and dosing decisions for norepinephrine ($vp_{2}$). The dosing decisions are formulated in Section 3.3. The reward function contains two objectives: maximizing the chance of survival and keeping vasopressors in dosage limit. Given a dataset $\mathcal{D}=\{(s_{t},a_{t},r_{t},s_{t+1})\}$ of historical ICU trajectories collected under behavioral policy $\pi_{b}$ (clinical practice), our goal is to learn an optimal policy $\pi^{*}$ that maximizes expected cumulative reward.
3.2 Q-Learning and conservative Q-Learning (CQL)
Standard Q-learning seeks to learn the action-value function $Q(s,a)$ representing the expected cumulative reward from taking action $a$ in state $s$ and following the optimal policy thereafter. The Q-function satisfies the Bellman optimality: $Q^{*}(s,a)=\mathcal{R}(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[\max_{a^{\prime}}Q^{*}(s^{\prime},a^{\prime})]$. In the offline setting, Q-learning minimizes the temporal difference (TD) mean-squared error:
$\mathcal{L}_{\text{TD}}(\theta)=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[\left(Q_{\theta}(s,a)-(r+\gamma\max_{a^{\prime}}Q_{\theta^{-}}(s^{\prime},a^{\prime}))\right)^{2}\right],$
where $Q_{\theta}$ is the learned Q-network with parameters $\theta$, and $Q_{\theta^{-}}$ is a target network updated periodically for stability. We apply a version of double Q-learning [^4] We use two Q nets, taking Q-value as $q=\text{min}(q_{1},q_{2})$; two target nets, taking target Q as $q^{\prime}=\text{min}(q_{1}^{\prime},q_{2}^{\prime})$. Standard Q-learning in offline settings may have over-estimation bias due to distributional shift, causing over-optimistic Q-value estimates.
As an improvement to Q-learning, CQL kumar2020conservative makes a correction by adding a conservative regularization term that penalizes Q-values for out-of-distribution (OOD) actions:
$\mathcal{L}_{\text{CQL}}(\theta)=\alpha\cdot\mathbb{E}_{s\sim\mathcal{D}}\left[\log\sum_{a}\exp(Q_{\theta}(s,a))-\mathbb{E}_{a\sim\mathcal{D}(a|s)}[Q_{\theta}(s,a)]\right]+\mathcal{L}_{\text{TD}}(\theta),$
where $\alpha$ controls the strength of the conservative penalty. When $\alpha=0$, it reduces to Q-learning.
3.3 Effective Dosing Action Space Design
We propose an action-space design to integrate clinical implementable practice with reinforcement learning. While continuous dosing variables are elegant, they may not agree with clinical realities of established protocols with discrete thresholds. Additionally, continuous action spaces compound RL research challenges due to exploration complexity and sparse rewards. Our interdisciplinary approach addresses this gap through three complementary action-space formulations: continuous dosage (for comparison), clinically meaningful discrete blocks, and directional stepwise adjustments. The action space $\mathcal{A}$={($vp_{1}$$vp_{2}$)} encompasses dual vasopressor dosing, enabling both computational tractability and clinical feasibility.
Continuous dosing (baseline). We use a continuous dosing variable for vasopressor norepinephrine allowing it to freely move move in a range. The action space is $\mathcal{A}=\{(vp_{1b},vp_{2c})\}$, where $vp_{1b}\in\{0,1\}$ is a binary variable, and $vp_{2c}\in[0,0.5]$ (mcg/kg/min).
Block discrete dosing. We define the dosing action to be the discrete values in the same range as the continuous dosing. The discrete values only vary across meaningful discrete dosing blocks. We explore different numbers of blocks to validate the impact of effective dosing. The action space is $\mathcal{A}=\{(vp_{1b},vp_{2d})\}$, and $vp_{2c}\in\text{range}(\delta/2,0.5,\delta)$ (mcg/kg/min) where $\delta$ is the minimum step [^5] Due to international clinical practice, we ensure the dosing is above zero for the first-line vasopressor..
Stepwise directional dosing. We define the the action space at each time step to be a directional and stepwise change. $\mathcal{A}=\{(vp_{1b},vp_{2s})\}$, and $vp_{2s}$ is a stepwise change from the current dose, for example, $vp_{2s}\in[{-0.1,-0.05,0,+0.05,+0.1}]$ (mcg/kg/min). The state space is extended to include a one-hot vector that stores and keeps track of the current block-discrete dosage of norepinephrine.
3.4 Recurrent Experience Replay with Effective Dosing
Our approach adapts the recurrent experience replay algorithm kapturowski2018recurrent to address the temporal dependencies in ICU vasopressor control. The algorithm employs a prioritized sequential buffer $\mathcal{RB}=\{\mathcal{S}_{e}\}$, storing treatment episodes $\mathcal{S}_{e}=\{(s_{t},a_{t},r_{t},s_{t+1})_{t\in e}\}$, to capture time-sensitive patterns of patient responses to vasopressor adjustments. In contrast to the sample-wise Q-function application in traditional Q-learning, we leverage an LSTM network to model the sequential nature of ICU states and interventions, recognizing that vasopressor effects unfold over time. The sum-tree data structure prioritizes episodes with larger temporal difference (TD) errors schaul2015prioritized , enabling learning on challenging clinical scenarios. This combination of recurrent processing and specialized action-space design addresses the ICU’s demands for both immediate responsiveness and patient stability.
4 Experiments
We apply the algorithms on the eICU pollard2018eicu , and MIMIC-IV johnson2023mimic datasets. The data is preprocessed with $16$ features and is split with ratios $70/15/15$ for train/val/test. In our RL formulation, the $vp_{1}$ (vasopressin) dosing is set as a binary variable because the clinical practice uses a single max dosage. We primarily experiment with the titration and weaning of $vp_{2}$ (norepinephrine)  [^6] See Appendix A and for detailed experiments. B. See Appendix C and D for model implementation and architecture. See Appendix E for hyperparameters and Appendix F for FQE off-policy evaluation.. The reward definition of the reinforcement learning model is summarized in Table 1.
Table 1: Definition of reward in the RL model.
Rule
Mortality +10 if the patient survived, and -10 otherwise
Vaso. penalty Each time step, aggregate to reward: $0.1\cdot(vp_{1}+vp_{2}\cdot 2-1.0))$
Our work compares several action-space formulations for vasopressor control in Table 2. We evaluate Q improvement ($\Delta Q$)—the gain in model-recommended actions over clinician actions—averaged across time steps. Each Q-function maps state–action pairs to a scalar value of expected future rewards, making $\Delta Q$ directly comparable across models. The Binary $vp_{1}$ model treats vasopressin as a binary variable. he Dual Mixed model represents $vp_{2}$ as a continuous variable in the range $[0,0.5]$ (mcg/kg/min). The Dual Block Discrete (BD) model discretizes $vp_{2}$ into $3$, 55, or $10$ bins. The Dual Stepwise model defines $vp_{2s}\in[{-\text{max\_step},...,-0.05,0,+0.05,...,+\text{max\_step}}]$ (mcg/kg/min). We report concordance (C.) between model recommendations and clinician actions. For $vp_{1}$, concordance is binary agreement; for $vp_{2}$, concordance holds when the recommended dose falls within the defined bin edges. Finally, the LSTM recurrent experience replay uses $vp_{2}$ discretized into 10 bins, trained with sequence length 5 and a 2-step burn-in. See model implemetation details in Appendix C.
Table 2: Comparison of RL models including Binary $vp_{1}$, Dual Mixed, Dual Block Discrete, Dual Stepwise models with the best conservatism levels ($\alpha$) selected with the validation set.
[…]
Clinician 38.8 0.000 100.0
Binary $vp_{1}$ 0.001 79.8 0.869 0.028 44.5
Dual Mixed 0.010 77.8 1.345 0.185 52.5
Dual BD. 3 bins 0.010 76.8 1.576 0.127 53.1 56.3
Dual BD. 5 bins 0.000 96.3 2.383 0.265 39.7 26.4
Dual BD 10 bins 0.000 92.3 4.673 0.309 42.4 13.3
Dual Stepwise max_step 0.1 0.0001 12.4 0.094 0.136 65.1 17.7
Dual Stepwise max_step 0.2 0.0000 97.3 4.021 0.511 38.1 11.7
LSTM BD 10 bins 0.0000 100.0 0.878 0.456 24.5
LSTM BD 10 bins 0.0001 100.0 0.866 0.398 20.4
We apply Fitted Q-Evaluation (FQE) off-policy evaluation (OPE) zhang2022off ; uehara2022review for all models in Table 3. For each model, we compute test set Q-values and fit a Gaussian to summarize their distribution. We compare model optimal policy with the clinician actions. The results show that block discrete models perform as well as continuous variable for $vp_{2}$. Stepwise directional with a larger max_step, has the largest 25.1% survivial improvement probability. Moreover, incorporating an LSTM with recurrent experience reply yields a survival improvement probability of 17.4%. This shows the benefits of capturing cross-time dependencies.
Table 3: FQE off-policy evaluation with the best $\alpha$ from validation. We define $\Delta$ Surv. Imp. Prob. (%) (Survival Improvement Probability) as the percentage of model recommended policy reaching higher reward value above the clinician mean reward minus 50% (PP(Model > Cli. Mean) - $50\%$), and $\Delta$ Q Mean (Mean Improvement) as the shift in the mean of fitted Gaussian for Model vs Clinician.
Model $\alpha$ Config $\Delta$ Surv. Imp. Prob. (%) $\Delta$ Q Mean Model M. Cli. M. Cohen’s d
Binary $vp_{1}$ 1e-3 0.8% 0.028 0.869 0.841 0.019
Dual Mixed 1e-2 6.3% 0.186 1.346 1.161 0.146
Dual BD 1e-2 3 bins 3.0% 0.127 1.576 1.450 0.073
Dual BD 0.0 5 bins 4.4% 0.265 2.383 2.119 0.106
Dual BD 0.0 10 bins 6.4% 0.309 4.673 4.364 0.151
Dual Stepwise 0.0 max_step 0.2 25.1% 0.511 4.021 3.510 0.503
LSTM BD 0.0 10 bins 13.9% 0.456 0.878 0.421 0.380
LSTM BD 1e-4 10 bins 17.4% 0.398 0.866 0.468 0.439
5 Conclusion and Future Work
This paper establishes action-space design as a critical bridge between reinforcement learning optimization and clinical implementation in ICU settings. By developing a novel action space, we convert theoretical optimal policies into actionable clinical protocols. Our proposed dual vasopressor control system for septic shock demonstrates that recurrent experience replay effectively captures the complex temporal dependencies inherent in ICU patient responses, achieving superior performance of over 15% survival improvement probability. These contributions move towards a blueprint for CDSS systems that clinicians could interpret and trust, and towards AI systems that not only compute optimal actions, but deliver them in forms that integrate into critical medical workflows.
Appendix A Fitted Q-Evaluation Illustrations
Figure 1: FQE comparison between learned patient model and clinician baseline (Binary $vp_{1}$).
Binary $vp_{1}$ Model: distribution of Q-values from model (red) and clinician (blue).
Figure 2: FQE comparison between the learned patient model and clinician baseline (Dual Mixed).
Dual Mixed Model: distribution of Q-values from model (red) and clinician (blue).
Figure 3: FQE comparison between the learned patient model and clinician baseline (Dual BD).
Block Discrete Model (10 bins): distribution of Q-values from model (red) and clinician (blue).
Figure 4: FQE comparison between the learned patient model and clinician baseline (Dual Stepwise).
Dual Stepwise Model (max_step=0.2): distribution of Q-values from model (red) and clinician (blue).
Figure 5: FQE comparison between the learned patient model and clinician baseline (LSTM BD).
LSTM BD (10 bins): distribution of Q-values from model (red) and clinician (blue).
Appendix B Detailed Experiment Results
Table 4: Comprehensive comparison of Q-learning models including Binary, Dual Mixed, and Dual Block Discrete (with 3, 5, 10 bins) variants across different conservatism levels ($\alpha$).
[…]
Clinician 38.8 0.000 0.000 100.0
Binary $vp_{1}$ 0.000 77.3 0.792 0.023 51.6
Dual Mixed 0.000 87.7 1.367 0.177 41.2
Dual Blk. Disc. 3 bins 0.000 71.4 1.734 0.123 55.2 57.1
Dual Blk. Disc. 5 bins 0.000 96.3 2.383 0.265 39.7 26.4
Dual Blk. Disc. 10 bins 0.000 92.3 4.673 0.309 42.4 13.3
Binary $vp_{1}$ 0.001 79.8 0.869 0.028 44.5
Dual Mixed 0.001 69.4 1.302 0.184 56.6
Dual Blk. Disc. 3 bins 0.001 82.2 1.614 0.110 46.9 55.5
Dual Blk. Disc. 5 bins 0.001 93.6 2.281 0.217 41.3 27.2
Dual Blk. Disc. 10 bins 0.001 96.6 4.445 0.292 39.1 14.1
Binary $vp_{1}$ 0.010 66.2 0.774 0.024 58.7
Dual Mixed 0.010 77.8 1.345 0.185 52.5
Dual Blk. Disc. 3 bins 0.010 76.8 1.576 0.127 53.1 56.3
Dual Blk. Disc. 5 bins 0.010 84.9 2.091 0.119 43.9 40.3
Dual Blk. Disc. 10 bins 0.010 92.8 4.072 0.245 41.4 16.6
Table 5: Comparison of Binary $vp_{1}$ and Dual Stepwise models with vasopressor persistence policy at different maximum step sizes.
[…]
Clinician 38.8 0.000 0.000 100.0
Binary $vp_{1}$ 0.001 79.8 0.869 0.028 44.5
Dual Mixed 0.010 77.8 1.345 0.185 52.5
max_step = 0.1 (mcg/kg/min)
Dual Stepwise 0.000000 34.6 0.046 0.103 61.1 17.9
Dual Stepwise 0.000100 12.4 0.094 0.136 65.1 17.7
Dual Stepwise 0.001000 82.9 0.252 0.128 47.4 24.5
Dual Stepwise 0.010000 45.9 -0.337 0.114 60.7 22.3
max_step = 0.2 (mcg/kg/min)
Dual Stepwise 0.000000 97.3 4.021 0.511 38.1 11.7
Dual Stepwise 0.000100 31.3 0.134 0.143 70.8 16.4
Dual Stepwise 0.001000 62.0 0.177 0.130 61.2 24.1
Dual Stepwise 0.010000 38.7 -0.184 0.254 64.6 13.5
Appendix C Model Implementation and Training
C.1 Double Q-learning
We use double-Q learning for all RL models in this paper. For our simplified version of the double-Q learning for stability, we adopt two Q-networks, and when Q-value is needed, the minimum scalar output across the two network is taken. Similarly, we adopt two target networks and take the minimum across them when computing the temporal difference (TD) error.
C.2 Binary vp1 Model
The Binary $vp_{1}$ model represents the simplest action space, treating vasopressin administration as a binary decision (on/off). The vp2 levels are used as part of the state. Training employs standard Q-learning TD loss with conservative logsumexp regularization over both action choices to prevent overestimation of out-of-distribution actions. The model processes states through fully connected networks with ReLU activations, using double Q-learning with target networks for stability.
C.3 Dual Mixed Model
The Dual Mixed model handles two action dimensions: $vp_{1}$ (binary, 0 or 1) and $vp_{2}$ (continuous, 0-0.5 mcg/kg/min). This continuous action space allows for fine-grained dosing control. For action selection of vp2, the model samples 50 candidate actions per state during both training and inference, evaluating Q-values for each to select optimal doses. The conservative Q-learning penalty is computed through importance sampling over randomly sampled actions from the continuous space.
C.4 Dual Block Discrete Model
The Dual Block Discrete model discretizes the continuous $vp_{2}$ action space into configurable bins while maintaining $vp_{1}$ as binary, creating a factored discrete action space with $2\times N$ (combined with the binary $vp_{1}$ action) total actions where $N$ is the number of $vp_{2}$ bins. The discretization uses uniform binning from 0 to 0.5 (mcg/kg/min) with bin centers used for continuous dose reconstruction. Q-values are computed for all discrete action combinations using one-hot encoding, enabling efficient batch evaluation. The conservative Q-learning penalty applies logsumexp over all valid discrete actions to maintain conservative value estimates.
C.5 Dual Stepwise Model
[…]
C.6 LSTM Block Discrete Model
The LSTM Block Discrete model extends the block discrete approach with sequential modeling capabilities. It processes patient trajectories through LSTM layers to capture temporal dependencies in treatment responses. The model uses sequence lengths of 5 timesteps with 2-step burn-in periods for hidden state initialization. Training employs a specialized replay buffer that maintains overlapping sequences with mortality-weighted sampling priorities. The architecture combines LSTM encoders (2 layers, 32 hidden units) with Q-networks that process concatenated LSTM outputs and state features. Actions are discretized into 10 norepinephrine dosing levels, with conservative Q-learning penalties computed over the discrete action space.
Appendix D Model Architectures
Table 6: Q Network Architecture (Binary/Mixed/Block/Stepwise)
Component Architecture Details
Input Layer State (state_dim) + Action (action_dim)
Hidden Layers [State, Action] $\xrightarrow{\text{FC (ReLU)}}$ 128 $\xrightarrow{\text{FC (ReLU)}}$ 128 $\xrightarrow{\text{FC (ReLU)}}$ 64
Output Layer 64 $\xrightarrow{\text{FC}}$ 1 (Q-value)
Initialization Xavier Uniform
Function $Q(s,a)\rightarrow\mathbb{R}$
Networks Dual Q-networks (Q1, Q2) with dual target networks
Table 7: LSTM Q Network Architecture (LSTM Block Discrete CQL)
Component Architecture Details
Input State (state_dim)
Feature Extractor [State] $\xrightarrow{\text{FC (ReLU)}}$ 32 $\xrightarrow{\text{Dropout(0.1)}}$
$\xrightarrow{\text{FC (ReLU)}}$ 32 $\xrightarrow{\text{Dropout(0.1)}}$
LSTM Layers 32 $\xrightarrow{\text{LSTM}}$ 2 layers $\times$ 32 units
(with dropout=0.1)
Output Layer (Q-head) 32 $\xrightarrow{\text{FC}}$ num_actions (Q-values)
Initialization Xavier Uniform
Function $Q(s,a)\rightarrow\mathbb{R}^{\text{num\_actions}}$
Networks Dual Q-networks with LSTM, dual target networks
Appendix E Hyperparameter Configuration
E.1 Shared Hyperparameters (All Models)
Table 8: Common Hyperparameters Across All Models
Hyperparameter Value Description
Learning Rate (lr) $10^{-3}$ Adam optimizer learning rate
Batch Size 128 Training batch size
Gamma ($\gamma$) 0.95 Discount factor for future rewards
Tau ($\tau$) 0.8 Soft target network update rate
Gradient Clipping 1.0 Maximum gradient norm
Epochs 100 Training epochs
Validation Batches 10 Batches for validation evaluation
Random Seed 42 For reproducibility
E.2 Model-Specific Configurations
Table 9: Model Configurations: Action Spaces and Hyperparameters
Action Space Key Parameters
Model Configuration Model Hyperparameters
Binary 1D binary (0/1) Binary
Dual Mixed 2D $vp_{1}$ binary, $vp_{2}$ continuous Dual Mixed num_samples: 50
$vp_{1}$: {0,1}, $vp_{2}$: [0,0.5] (continuous $vp_{2}$ selection)
[…]
$vp_{1}$: binary, $vp_{2}$: N bins
[…]
$vp_{1}$: binary, $vp_{2}$: M steps step_size: 0.05
LSTM Block N discrete ($vp_{2}$ only) LSTM Block seq_len: 5, burn_in: 2
Appendix F Offline Policy Evaluation with Fitted Q-Evaluation (FQE)
To evaluate the learned policies against historical clinician decisions without online deployment, we employed Fitted Q-Evaluation (FQE) with Gaussian distribution modeling. This approach computes Q-values for both the learned policy and clinician actions on the held-out test set, enabling direct comparison of expected returns. For each trained model, we extract Q-values by evaluating state-action pairs from the test trajectories where actions are either from the learned policy (model Q-values) or from historical clinician decisions (clinician Q-values). The resulting Q-value distributions are then fitted with Gaussian distributions using maximum likelihood estimation, providing parametric representations characterized by mean ($\mu$) and standard deviation ($\sigma$) for both the model and clinician policies. This parametric approach enables computation of key metrics including the mean improvement gap (difference in expected returns), probability of improvement (likelihood that model Q-values exceed clinician Q-values), and effect size (Cohen’s d) to quantify the magnitude of policy differences.
The FQE analysis also generates visualizations including overlapping histograms with fitted Gaussian curves, cumulative distribution functions (CDFs) showing the probability mass distribution of Q-values, and improvement probability curves that illustrate the likelihood of the model outperforming various Q-value thresholds. The probability of improvement at the clinician’s mean Q-value serves as a critical metric, indicating how often the learned policy is expected to achieve better outcomes than historical treatment decisions. Additionally, the analysis computes Cohen’s d as a standardized effect size measure. The larger the value of Cohen’s d, the more meaningful are the difference between the means of two groups. This framework provides quantitative evidence for policy improvement while accounting for the inherent uncertainty in Q-value estimates. The FQE analysis was performed for policy performance across different model architectures and hyperparameter configurations. The FQE numerical results are given in Table 3. The histogram and sigmoid CDF visualizations are shown in Appendix A.